The length of the rectangle is 1.5 times its width and the perimeter is 75 feet
I did guess and check but i couldnt find the answer

Answer by texttutoring(324)

You have 2 variables (Length and Width). You need to come up with 2 equations.
the perimeter is 75 feet. A rectangle is composed of 2 lengths and 2 widths.
Your equations are:
Eqn 1: 2w+2L=75
Eqn 2: 1.5w=L
Substitute Eqn 2 into Eqn 1:
2w+2(1.5w)=75
2w+3w=75
5w=75
w=15
Use Eqn 2 to find L:
1.5w=L
1.5*15=L
L=22.5
The width is 15 feet, the length is 22.5 feet.